OpenStudy (anonymous):

how do you solve (3/x-1)+(2/x)=4?

OpenStudy (anonymous):

How would you solve*

OpenStudy (anonymous):

look

OpenStudy (anonymous):

\[\frac{ 3 }{ x }-1+\frac{ 2 }{ x }=4\] \[\frac{ 3 }{ x }+\frac{ 2 }{ x }=5\] \[\frac{ 5 }{ x }=5\] 5x=5 "divided by 5" x=1

OpenStudy (anonymous):

no 3/(x-1)+2/x=4

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

is that the real question or what ??

OpenStudy (anonymous):

yes. Im sorry the real question is 3/(x-1)+2/x=4

OpenStudy (anonymous):

it's okay

OpenStudy (anonymous):

give me a min while writing it